OpenStudy (anonymous):

Using the Pythagorean theorem, find the measure of the hypotenuse when a = 2.4in and b = 6.4in.

OpenStudy (jackellyn):

\[a ^{2}+b ^{2}=c ^{2}\] Plug in the numbers (2.4)^2 + (6.4)^2 =c^2 46.72=c^2 square root both sides \[\sqrt{46.72}=\sqrt{c ^{2}}\] c = 6.835203